OpenStudy (anonymous):

Two trains leave stations 560 miles apart at the same time and travel toward each other. One train travels at 95 miles per hour while the other travels at 105 miles per hour. How long will it take for the two trains to meet?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

The answer is: 2.8 hours

OpenStudy (aaronq):

 train 1 (x=0) \(x_f=x_i+v\Delta t=0+(95mi/hr)\Delta t\) train 1 (x=0) \(x_f=x_i+v\Delta t=560~mi+(-105mi/hr)\Delta t\) same final position, \(x_f\). We have equation: \((95mi/hr)\Delta t=560~mi-(105mi/hr)\Delta t\) Now we just solve for \(\Delta t\)

OpenStudy (aaronq):

i made a mistake (copying and pasting my own work) the second time it says "train 1 (x=0)" was supposed to say "train 2 (x=560)"